1. Point charges q1 = −4.5 nC and q2 = +4.5 nC are separated by 3.1 mm, forming an electric

dipole. (a) Find the electric dipole moment (magnitude and direction). (b) The charges are in a

uniform electric field whose direction makes an angle of 36.9 ̊ with the line connecting the charges.

What is the magnitude of this field if the torque exerted on the dipole has a magnitude 7.2 × 10−9N ∙

m?


1
Expert's answer
2022-01-10T09:16:07-0500

Answer

Charge

"| q_1|=|q_2|=4.5nC."

Distance d=3.1mm

a) electric dipole

"P=qd\\=4.5nC\times3.1mm\\=13.95\times10^{-12}C-m"

Direction is charge q1 to q2.


b) now electric field is given

"E=\frac{\tau }{P\sin\theta}"

Putting all values

"E=\frac{7.2\times10^{-9}}{13.95\times10^{-12}\sin36.9°}"

"=\frac{7200}{13.95\times.596}"

"=867.85V\/m"